# Computational lexical analysis of Flamenco genres

### v1

@@ III Discussion
In tangos, there are characteristic mentions of geographical locations such as the town of Utrera or the Seville neighborhood of Triana, including its famous puente ’bridge’. Dios ’God’ is also present, surrounded by other religious references such as Virgen ’Virgin’ or Humildad ’Humility’ in reference to the Christ of Humility. Additionally, around guerra ’war’ we find allusions to historical conflicts, specifically the Cuban and French wars, are viewed through a Spanish historical lens.
@@
Overall, tangos engage deeply with emotions, covering a spectrum from sorrow (pena ’pain’, fatigas ’fatigues, pains’), to love, happiness and joy (querer ’to want/love’, bonita ’beautiful (fem. sing.)’, alegría ’joy’).
@@
Finally, tientos share much of their semantic content with tangos, encompassing human emotions and religious themes. This is observed in the use of lexicon such as querer ’to want/love’, pena ’pain, sorrow’, Dios ’God’, Virgen ’Virgin’, and guerra ’war’. Nonetheless, references to geographical places are less frequent. Predominantly, the essential vocabulary of tientos pertains to formulaic expressions, such as serrana ’mountain woman’, vente ’you come’, conmigo ’with me’, daré ’I will give’ in
@@
’Come with me, mountain girl
@@
Vocabulary not only characterizes each Flamenco genre but also establishes relationships among the styles. In Fig. 13 we observe four distinct branches converging at bulerías, serving as a unifying node. These branches encompass a gipsy branch involving seguiriyas and soleá, a branch (presumably) originating from Málaga comprising malagueñas and fandangos, and the branch of tangos, encompassing tangos and tientos. Importantly, the MST representation effectively encapsulates known historical relationships among palos. For instance, our result aligns with established historical accounts such as the derivation of tientos from a deceleration of tangos ([14], p. 338), and sheds light on lesser-known connections, such as the relationship between malagueñas and fandangos. While the origin of the latter is discussed by various theories ([14], p. 420), our finding supports the understanding of current fandangos as traditional songs originating from Málaga, thereby linking them to malagueñas [43].
@@ IV Conclusions
First, we have addressed RQ1, can Flamenco palos be classified solely by means of their lexicon? If so, what is the characteristic lexicon of each palo? Our findings demonstrate that, with varying degrees of accuracy, it is indeed possible to automatically classify Flamenco songs into their respective palos based solely on their lexicon. This discovery holds significant relevance for Flamenco palo classification, which have thus far relied largely on subjective judgment.
@@
Finally, our study is not without its limitations. Unlike mainstream genres like reggaeton, pop, or rock, Flamenco’s limited availability of large lyric databases complicates analysis, further compounded by its rich collection of styles. While our study benefited from a reasonably curated corpus, the dataset still lacked representation of numerous palos. Moreover, among those genres that were sufficiently represented, there were notable disparities in representation: some were significantly more represented than others, underscoring the corpus-dependency of our analysis and conclusions. Future research should aim to address these limitations by expanding the scope of data collection efforts, thereby enhancing the robustness and generalizability of findings. Moreover, a change in the preprocessing stage, stop words, or other parameters of our analysis may yield different results, but we believe that the main conclusions of the current study will remain valid. Also, incorporating advanced text processing techniques such as lemmatization or stemming could refine our analysis [44] and yield deeper insights into the lexical nuances of Flamenco lyrics. As Flamenco continues to evolve, marked by contemporary references like Moneo’s mention of the coronavirus in recent performances [45], ongoing research efforts must adapt to capture and interpret these dynamic shifts in the genre’s lexical content. To sum up, embracing quantitative methods and interdisciplinary applications have allowed us to uncover Flamenco’s linguistic and cultural legacy, thereby deepening our appreciation for this iconic musical tradition.
@@ Acknowledgements.
The authors thank Norman P. Kliman for useful discussions about the dataset [23]. This work was partially supported by the Spanish State Research Agency (MICIU/AEI/10.13039/501100011033) and FEDER (UE) under project APASOS (PID2021-122256NB-C21) and the María de Maeztu project CEX2021-001164-M, and by the Government of the Balearic Islands CAIB fund ITS2017-006 under project CAFECONMIEL (PDR2020/51).
@@ A.1 Empirical laws
We check its adherence to both Zipf’s and Heaps’ laws, as illustrated in Figs. 14 and 15, respectively. The observed exponents agree with many other corpora widely employed in the literature [46].
Figure 14: Log-log scale distribution of tokens in the entire and unfiltered corpus, fitted to Zipf’s law, which relates the number of occurrences of a word with its rank as a power law with an exponent of $-1.06$ (dashed line).
Figure 15: Log-log scale of the relationship between corpus size, $L$, and vocabulary size, $|V|$, adhering to Heaps’ law. We plot the relationship for each genre independently (green lines) and for the global corpus (thick black line), in both cases for unfiltered texts. The sublinear scaling observed underscores the increasing lexical diversity as the corpus grows, with an exponent of $0.55$ for large amounts of tokens (dashed line).
@@ B.1 Multinomial Naive Bayes
[…]
@@
$\text{Score}(C|d)$ $=\log[P(C)]$
$+\sum_{w\in d}\left(\log\left[P(w|C)\right]\text{TF-IDF}(w,d)\right),$ (4)
@@
With these values of $\alpha$ and training size, the model behaves as shown in the confusion matrix of Fig. 5 and in the accuracy ranking of Fig. 6. The accuracies are reasonably good for most of the styles. A few styles do not perform as well, due to lexical similarities which are discussed in Section II.
@@ B.2 Essential words
B.3 Distance and clustering
@@ C.1 Alegrías
In Fig. 8a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 9 we show the complete $\log P\left(w|C\right)$ distribution for alegrías.
The essential words for alegrías in descending order of $P\left(w|C\right)$ are the following:
Cádiz ’Cadiz’, llevar ’carry’, Cuando ’When’, dirá ’will say’, conmigo ’with me’, da ’gives’, mandilón ’submissive, pussy-whipped’, veo ’I see’, llaman ’they call’, Muralla Real ’Royal Wall’, vayas ’you go’, ’sea’, voy ’I go’, mar, adónde ’where to’, vengas ’you come’, más ’more’, quiero ’I want’, dar ’to give’, Que/Qué ’Which/What’, bonita ’beautiful (fem.)’, contigo ’with you’, nombre ’name’, darte ’to give you’, vas ’you go’, niña ’girl’, pasando ’happening’, paseíto ’little walk’, vueltecita ’little turn’, tacón ’heel’, pena ’pain’, patrona ’patroness’, saber ’to know’, sale ’it goes out’, Zaragoza ’Zaragoza’, alegria ’joy’, doble ’double’, relicario ’reliquary’, entender ’to understand’, Navarra ’Navarra’, taconear ’to tap (with heels)’, bota ’boot’, Rosario ’Rosary’, arena ’sand’, San Fernando ’Saint Fernando’, mientan ’they lie’, buenos ’good (masc. pl.)’, barrio ’neighborhood’, tiro ’throw’, están ’they are’, duro ’hard’, madre ’mother’, entrada ’entrance’, valgo ’I am worth’, señora ’lady’, viva ’long life to’, muero ’I die’, titirimundis ’puppets’, dos ’two’, barquero ’boatman’, confianza ’trust’, mujeres ’women’, alma ’soul’, camino ’road’, plata ’silver’, nadie ’nobody’, Una ’One (fem.)’, pago ’I pay’, verte ’to see you’, cayó ’it fell’, prima ’cousin (fem.)’, Yo ’I’, Me ’Me’.
@@ C.2 Bulerías
In Fig. 17 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 18 we show the complete $\log P\left(w|C\right)$ distribution for bulerías.
Figure 17: The 15 words with highest $\log P\left(w|C\right)$ for bulerías.
Figure 18: Complete $\log P\left(w|C\right)$ distribution for bulerías over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for bulerías in descending order of $P\left(w|C\right)$ are the following:
voy ’I go’, quiero ’I want/love’, más ’more’, madre ’mother’, gitana ’gypsy (fem. sing.)’, Yo ’I’, usted ’you (formal)’, Que/Qué ’Which/What’, Dios ’God’, noche ’night’, mía ’mine (fem. sing.)’, va ’goes’, Ay, tan ’so (comp.)’, amor ’love’, vino ’wine’, van ’they go’, nadie ’nobody’, vida ’life’, cara ’face’, alma ’soul’, ahora ’now’, contigo ’with you’, gente ’people’, querer ’to love’, Esta ’This (fem.)’, gitano ’gypsy (masc. sing.)’, ver ’to see’, niña ’girl’, mío ’mine’, quiera ’he/she wants’, dos ’two’, calle ’street’, agua ’water’, pena ’pain, sorrow’, Por ’For’, dinero ’money’, corazón ’heart’, pelo ’hair’, quieres ’you want/love’, día ’day’, conmigo ’with me’, vaya ’he/she goes’, Tu ’You’, quiere ’he/she wants’, ser ’to be’, vas ’you go’, casa ’house’, tenía ’he/she had’, aunque ’although’, lleva ’he/she wears’, llevo ’I wear’, mal ’bad’, río ’river’, padre ’father’, digo ’I say’, puedo ’I can’, perdido ’lost’, ojos ’eyes’, tres ’three’, será ’it will be’, No ’No’, mañana ’morning’, mama ’mom’, caballo ’horse’, Cuando ’When’, vera ’side, proximity’, viene ’he/she comes’, grande ’big’, bonita ’beautiful (fem. sing.)’, flores ’flowers’, Él ’He’, hacer ’to do’, cuatro ’four’, quisiera ’I would like’, Virgen ’Virgin’, mundo ’world’, alegría ’joy’, Me ’Me’, niño ’child (masc.)’, mala ’bad’, aquél ’that one’, puerta ’door’, comer ’to eat’, prima ’cousin (fem.)’, campanas ’bells’, cuerpo ’body’, camino ’path’, paso ’step’, dicen ’they say’, loca ’crazy’, Sevilla ’Seville’, María ’María’, Amparo ’Amparo’, dice ’he/she says’, mano ’hand’, daba ’he/she gave’, primo ’cousin (masc.)’.
@@ C.3 Fandangos
In Fig. 19 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 20 we show the complete $\log P\left(w|C\right)$ distribution for fandangos.
Figure 19: The 15 words with highest $\log P\left(w|C\right)$ for fandangos.
Figure 20: Complete $\log P\left(w|C\right)$ distribution for fandangos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for fandangos in descending order of $\log P\left(w|C\right)$ in descending order of $P\left(w|C\right)$ are the following:
mujer ’woman’, día ’day’, madre ’mother’, más ’more’, quería ’I wanted/I loved’, querer ’to want/love’, llorando ’crying’, vi ’I saw’, hombre ’man’, vida ’life’, quiero ’I want/love’, Dios ’God’, tan ’so (comp.)’, Yo ’I’, mía ’mine (fem. sing.)’, voy ’I go’, mundo ’world’, pena ’pain, sorrow’, noche ’night’, ser ’to be’, nadie ’nobody’, quieres ’you want/love’, llorar ’to cry’, De ’Of’, había ’there was’, ver ’to see’, vivir ’to live’, cruz ’cross’, cara ’face’, alma ’soul’, solo ’alone’, dije ’I said’, también ’also’, corazón ’heart’, tenía ’he/she had’, flores ’flowers’, tener ’to have’, agua ’water’, Él ’He’, nunca ’never’, contigo ’with you’, llorado ’cried’, mira ’look’, quiere ’he/she wants’, cinco ’five’, querido ’dear, loved’, vez ’time’, mío ’mine’, loco ’crazy’, penas ’pains, sorrows’, ahora ’now’, pecho ’chest’, dos ’two’, pensamiento ’thought’, paró ’stopped’, traición ’betrayal’, llegar ’to arrive’, boca ’mouth’, gitana ’gypsy (fem.)’, Con ’With’, quiera ’he/she wants’, muerta ’dead (fem. sing.)’, verdad ’truth’, niños ’little boys’, puedo ’I can’, puerta ’door’, Que/Qué ’Which/What’, morena ’brunette’, No ’No’, dinero ’money’, lloran ’they cry’, vera ’side, proximity’, trigal ’wheat field’, sangre ’blood’, Vela ’Candle’, pienso ’I think’, muere ’he/she dies’, acababa ’it was ending’, ninguna ’none (fem.)’, morir ’to die’, días ’days’, decía ’he/she said’, Porque ’Because’, va ’he/she goes’, tarde ’late’, puedes ’you can’, clara ’clear (fem.)’, manos ’hands’, contado ’counted’, sentidos ’senses’, cariño ’affection’, quieras ’you want’, cambio ’change’, gente ’people’.

### v2

@@ III Discussion
In tangos, there are characteristic mentions of geographical locations such as the town of Utrera or the Seville neighborhood of Triana, including its famous puente ‘bridge’. Dios ‘God’ is also present, surrounded by other religious references such as Virgen ‘Virgin’ or Humildad ‘Humility’ in reference to the Christ of Humility. Additionally, around guerra ‘war’ we find allusions to historical conflicts, specifically the Cuban and French wars, are viewed through a Spanish historical lens.
@@
Overall, tangos engage deeply with emotions, covering a spectrum from sorrow (pena ‘pain’, fatigas ‘fatigues, pains’), to love, happiness and joy (querer ‘to want/love’, bonita ‘beautiful (fem. sing.)’, alegría ‘joy’).
@@
Finally, tientos share much of their semantic content with tangos, encompassing human emotions and religious themes. This is observed in the use of lexicon such as querer ‘to want/love’, pena ‘pain, sorrow’, Dios ‘God’, Virgen ‘Virgin’, and guerra ‘war’. Nonetheless, references to geographical places are less frequent. Predominantly, the essential vocabulary of tientos pertains to formulaic expressions, such as serrana ‘mountain woman’, vente ‘you come’, conmigo ‘with me’, daré ‘I will give’ in
@@
‘Come with me, mountain girl
@@
Vocabulary not only characterizes each Flamenco genre but also establishes relationships among the styles. In Fig. 10(b) we observe four distinct branches converging at bulerías, serving as a unifying node. These branches encompass a gipsy branch involving seguiriyas and soleá, a branch (presumably) originating from Málaga comprising malagueñas and fandangos, and the branch of tangos, encompassing tangos and tientos. Importantly, the MST representation effectively encapsulates known historical relationships among palos. For instance, our result aligns with established historical accounts such as the derivation of tientos from a deceleration of tangos ([18], p. 338), and sheds light on lesser-known connections, such as the relationship between malagueñas and fandangos. While the origin of the latter is discussed by various theories ([18], p. 420), our finding supports the understanding of current fandangos as traditional songs originating from Málaga, thereby linking them to malagueñas [53].
@@ IV Conclusions
First, we have addressed RQ1, Can Flamenco songs be properly assigned to their corresponding palo solely by means of their lexicon? If so, what is the characteristic lexicon of each palo? Our findings demonstrate that, with varying degrees of accuracy, it is indeed possible to automatically classify Flamenco songs into their respective palos based solely on their lexicon. This discovery holds significant relevance for Flamenco palo classification, which has traditionally relied on subjective assignments by authors or record labels, guided by stylistic conventions. Our findings quantitatively demonstrate that these conventions are also reflected in consistent lexical patterns.
@@
Finally, our study is not without its limitations. Unlike mainstream genres like reggaeton, pop, or rock, Flamenco’s limited availability of large lyric databases complicates analysis, further compounded by its rich collection of styles. While our study benefited from a reasonably curated corpus, the dataset still lacked representation of numerous palos, meaning our conclusions are only reliable for the eight main palos analyzed. Furthermore, within the genres that were sufficiently represented, there were significant disparities in representation, with some palos more heavily represented than others. This highlights the corpus-dependence of our analysis and conclusions. Future research should address these limitations by expanding data collection efforts, which would strengthen the robustness and generalizability of the findings. With a more comprehensive dataset, more advanced techniques such as BERT, Word2Vec, or Convolutional and Recurrent Neural Networks, could be employed. For example, training an sBERT model on Flamenco lyrics could capture word relationships at a deeper level than our machine learning approach. Currently, Spanish language models do not fully capture the regional variants often found in Flamenco lyrics, which often differ from the standard Spanish found in typical training datasets. We are hopeful that PlanTL-GOB-ES/roberta-base-bne [12] and BSC-LT/mRoBERTa [29] models will address this gap. Future work should focus on applying these advanced technologies with properly trained datasets.
While our unigram approach has already uncovered complex formulaic structures, further work could explore $n$-grams with $n\geq 2$. This would capture both semantic and syntactic relationships that extend beyond the individual meanings of unigrams, and would allow for the automation of the concatenation process shown in Sec. A.2.
Moreover, a change in the preprocessing stage, stop words, or other parameters of our analysis may yield different results, but we believe that the main conclusions of the current study will remain valid. Also, incorporating advanced text processing techniques such as lemmatization or stemming could refine our analysis [6] and yield deeper insights into the lexical nuances of Flamenco lyrics. As Flamenco continues to evolve, marked by contemporary references like Moneo’s mention of the coronavirus in recent performances [35], ongoing research efforts must adapt to capture and interpret these dynamic shifts in the genre’s lexical content, therefore allowing for a diachronic analysis of the lexicon used in Flamenco lyrics. To sum up, embracing quantitative methods and interdisciplinary applications have allowed us to uncover Flamenco’s linguistic and cultural legacy, thereby deepening our appreciation for this iconic musical tradition.
@@ Acknowledgements.
The authors thank Norman P. Kliman, José María Velázquez-Gaztelu and Joaquín Guindo for useful discussions and references. This work was partially supported by the Spanish State Research Agency (MICIU/AEI/10.13039/501100011033) and FEDER (UE) under project APASOS (PID2021-122256NB-C21) and the María de Maeztu project CEX2021-001164-M, and by the Government of the Balearic Islands CAIB fund ITS2017-006 under project CAFECONMIEL (PDR2020/51).
@@ A.1 Empirical laws
We check its adherence to both Zipf’s and Heaps’ laws, as illustrated in Fig. 11. The observed exponents agree with many other corpora widely employed in the literature [3, 49].
Figure 11: (a) Log-log scale distribution of tokens in the entire and unfiltered corpus, fitted to Zipf’s law, which relates the number of occurrences of a word with its rank as a power law with an exponent of 1.06-1.06 (dashed line), and (b) log-log scale of the relationship between corpus size, $L$, and vocabulary size, $|V|$, adhering to Heaps’ law. We plot the relationship for each genre independently (green lines) and for the global corpus (thick black line), in both cases for unfiltered texts and for a given disordered realization. The sublinear scaling observed underscores the increasing lexical diversity as the corpus grows, with an exponent of $0.55$ for large amounts of tokens (dashed line).
@@ B.1 Multinomial Naive Bayes
$\text{TF-IDF}(w,d)=\bar{tf}(w,d)\left[1+\log\left(\frac{|D|}{df(w)}\right)\right],$ (2)
@@
$\text{Score}(C|d)=\log[P(C)]+\sum_{w\in d}\left(\log\left[P(w|C)\right]\text{TF-IDF}(w,d)\right),$ (4)
@@
With these values of $\alpha$ and training size, the model behaves as shown in the confusion matrix of Fig. 4 and in the accuracy ranking of Fig. 5(a). The accuracies are reasonably good for most of the styles. A few styles do not perform as well, due to lexical similarities which are discussed in Section II.
@@ B.2 Essential words
B.3 Hapax legomena
Additionally, let us examine the distribution of the ratio of palo-hapax legomena, i.e., the set of words that appear only in one palo and not in the others, $r_{\mathrm{h}}$, for each song of each palo,
$r_{\mathrm{h}}=\frac{\text{\# {palo}-hapax}}{|V_{\mathrm{s}}|},$ (5)
where $|V_{\mathrm{s}}|$ is the size of the vocabulary of a song.
In Fig. 13 we show the distribution of $r_{\mathrm{h}}$ for the filtered dataset. While distinct patterns may be observed, the distributions across all palos are generally quite similar. Overall, all distributions have their peaks located around $r_{\mathrm{h}}=0.1$, which is lower than $r_{\mathrm{h}}=0.5$, the pattern commonly observed in large corpora for corpus-level hapax legomena [5, 43].
Figure 13: Kernel density estimation [11] of the distribution of hapax legomena found in the filtered lyrics of each palo, following the acronyms of Table 1.
Hapax legomena at palo level are considered as relevant in certain ocasions by the MNB classifier fed with TF-IDF scores. Indeed, if we compare the palo-unique lexicon with the essential words computed in our manuscript, we find several coincidences, as seen in Table 4.
palo Genre-unique Essential Shared
A 376 73 5
B 2288 98 0
F 747 94 1
M 135 15 1
Se 379 172 5
So 746 182 0
Ta 624 57 0
Ti 273 55 0
Table 4: Amount of genre-unique lexicon and essential words for each palo, including shared word totals.
B.4 Distance and clustering
@@ C.1 Alegrías
In Fig. 6a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 7 we show the complete $\log P\left(w|C\right)$ distribution for alegrías. The essential words for alegrías in descending order of $P\left(w|C\right)$ are the following:
Cádiz ‘Cadiz’, llevar ‘carry’, Cuando ‘When’, dirá ‘will say’, conmigo ‘with me’, da ‘gives’, mandilón ‘submissive, pussy-whipped’, veo ‘I see’, llaman ‘they call’, Muralla Real ‘Royal Wall’, vayas ‘you go’, ’sea’, voy ‘I go’, mar, adónde ‘where to’, vengas ‘you come’, más ‘more’, quiero ‘I want’, dar ‘to give’, Que/Qué ‘Which/What’, bonita ‘beautiful (fem.)’, contigo ‘with you’, nombre ‘name’, darte ‘to give you’, vas ‘you go’, niña ‘girl’, pasando ‘happening’, paseíto ‘little walk’, vueltecita ‘little turn’, tacón ‘heel’, pena ‘pain’, patrona ‘patroness’, saber ‘to know’, sale ‘it goes out’, Zaragoza ‘Zaragoza’, alegria ‘joy’, doble ‘double’, relicario ‘reliquary’, entender ‘to understand’, Navarra ‘Navarra’, taconear ‘to tap (with heels)’, bota ‘boot’, Rosario ‘Rosary’, arena ‘sand’, San Fernando ‘Saint Fernando’, mientan ‘they lie’, buenos ‘good (masc. pl.)’, barrio ‘neighborhood’, tiro ‘throw’, están ‘they are’, duro ‘hard’, madre ‘mother’, entrada ‘entrance’, valgo ‘I am worth’, señora ‘lady’, viva ‘long life to’, muero ‘I die’, titirimundis ‘puppets’, dos ‘two’, barquero ‘boatman’, confianza ‘trust’, mujeres ‘women’, alma ‘soul’, camino ‘road’, plata ‘silver’, nadie ‘nobody’, Una ‘One (fem.)’, pago ‘I pay’, verte ‘to see you’, cayó ‘it fell’, prima ‘cousin (fem.)’, Yo ‘I’, Me ‘Me’.
@@ C.2 Bulerías
In Fig. 14(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 14(b) we show the complete $\log P\left(w|C\right)$ distribution for bulerías. The essential words for bulerías in descending order of $P\left(w|C\right)$ are the following:
Figure 14: (a) The 15 words with highest $\log P\left(w|C\right)$ for bulerías, and (b) complete $\log P\left(w|C\right)$ distribution for bulerías over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
voy ‘I go’, quiero ‘I want/love’, más ‘more’, madre ‘mother’, gitana ‘gypsy (fem. sing.)’, Yo ‘I’, usted ‘you (formal)’, Que/Qué ‘Which/What’, Dios ‘God’, noche ‘night’, mía ‘mine (fem. sing.)’, va ‘goes’, Ay, tan ‘so (comp.)’, amor ‘love’, vino ‘wine’, van ‘they go’, nadie ‘nobody’, vida ‘life’, cara ‘face’, alma ‘soul’, ahora ‘now’, contigo ‘with you’, gente ‘people’, querer ‘to love’, Esta ‘This (fem.)’, gitano ‘gypsy (masc. sing.)’, ver ‘to see’, niña ‘girl’, mío ‘mine’, quiera ‘he/she wants’, dos ‘two’, calle ‘street’, agua ‘water’, pena ‘pain, sorrow’, Por ‘For’, dinero ‘money’, corazón ‘heart’, pelo ‘hair’, quieres ‘you want/love’, día ‘day’, conmigo ‘with me’, vaya ‘he/she goes’, Tu ‘You’, quiere ‘he/she wants’, ser ‘to be’, vas ‘you go’, casa ‘house’, tenía ‘he/she had’, aunque ‘although’, lleva ‘he/she wears’, llevo ‘I wear’, mal ‘bad’, río ‘river’, padre ‘father’, digo ‘I say’, puedo ‘I can’, perdido ‘lost’, ojos ‘eyes’, tres ‘three’, será ‘it will be’, No ‘No’, mañana ‘morning’, mama ‘mom’, caballo ‘horse’, Cuando ‘When’, vera ‘side, proximity’, viene ‘he/she comes’, grande ‘big’, bonita ‘beautiful (fem. sing.)’, flores ‘flowers’, Él ‘He’, hacer ‘to do’, cuatro ‘four’, quisiera ‘I would like’, Virgen ‘Virgin’, mundo ‘world’, alegría ‘joy’, Me ‘Me’, niño ‘child (masc.)’, mala ‘bad’, aquél ‘that one’, puerta ‘door’, comer ‘to eat’, prima ‘cousin (fem.)’, campanas ‘bells’, cuerpo ‘body’, camino ‘path’, paso ‘step’, dicen ‘they say’, loca ‘crazy’, Sevilla ‘Seville’, María ‘María’, Amparo ‘Amparo’, dice ‘he/she says’, mano ‘hand’, daba ‘he/she gave’, primo ‘cousin (masc.)’.
@@ C.3 Fandangos
In Fig. 15(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 15(b) we show the complete $\log P\left(w|C\right)$ distribution for fandangos. The essential words for fandangos in descending order of $\log P\left(w|C\right)$ in descending order of $P\left(w|C\right)$ are the following:
Figure 15: (a) The 15 words with highest $\log P\left(w|C\right)$ for fandangos, and (b) complete $\log P\left(w|C\right)$ distribution for fandangos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
mujer ‘woman’, día ‘day’, madre ‘mother’, más ‘more’, quería ‘I wanted/I loved’, querer ‘to want/love’, llorando ‘crying’, vi ‘I saw’, hombre ‘man’, vida ‘life’, quiero ‘I want/love’, Dios ‘God’, tan ‘so (comp.)’, Yo ‘I’, mía ‘mine (fem. sing.)’, voy ‘I go’, mundo ‘world’, pena ‘pain, sorrow’, noche ‘night’, ser ‘to be’, nadie ‘nobody’, quieres ‘you want/love’, llorar ‘to cry’, De ‘Of’, había ‘there was’, ver ‘to see’, vivir ‘to live’, cruz ‘cross’, cara ‘face’, alma ‘soul’, solo ‘alone’, dije ‘I said’, también ‘also’, corazón ‘heart’, tenía ‘he/she had’, flores ‘flowers’, tener ‘to have’, agua ‘water’, Él ‘He’, nunca ‘never’, contigo ‘with you’, llorado ‘cried’, mira ‘look’, quiere ‘he/she wants’, cinco ‘five’, querido ‘dear, loved’, vez ‘time’, mío ‘mine’, loco ‘crazy’, penas ‘pains, sorrows’, ahora ‘now’, pecho ‘chest’, dos ‘two’, pensamiento ‘thought’, paró ‘stopped’, traición ‘betrayal’, llegar ‘to arrive’, boca ‘mouth’, gitana ‘gypsy (fem.)’, Con ‘With’, quiera ‘he/she wants’, muerta ‘dead (fem. sing.)’, verdad ‘truth’, niños ‘little boys’, puedo ‘I can’, puerta ‘door’, Que/Qué ‘Which/What’, morena ‘brunette’, No ‘No’, dinero ‘money’, lloran ‘they cry’, vera ‘side, proximity’, trigal ‘wheat field’, sangre ‘blood’, Vela ‘Candle’, pienso ‘I think’, muere ‘he/she dies’, acababa ‘it was ending’, ninguna ‘none (fem.)’, morir ‘to die’, días ‘days’, decía ‘he/she said’, Porque ‘Because’, va ‘he/she goes’, tarde ‘late’, puedes ‘you can’, clara ‘clear (fem.)’, manos ‘hands’, contado ‘counted’, sentidos ‘senses’, cariño ‘affection’, quieras ‘you want’, cambio ‘change’, gente ‘people’.